OpenStudy (anonymous):

Find the radius or diameter of each circle with the given dimensions. d= 4 ft

OpenStudy (anonymous):

Reverse the equation D=2*r and you get D/2=r so 4/2 = 2 ft

OpenStudy (anonymous):

r=d/2 r=4/2 r = 2 ft

OpenStudy (anonymous):

ok